OpenStudy (anonymous):

Please, Help!! The greatest and the least resultant of 2 forces acting at a point is 10Newton and 6Newton respectively. If each force is increased by 3Newton, Find the resultant of new forces when acting at a point at an angle of 90Degree with each other.

OpenStudy (anonymous):

@ganeshie8 @.Sam. @jim_thompson5910 @LoveYou*69 @Preetha @John_ES @Mertsj @UsukiDoll @soty2013 @SnuggieLad @ryan123345 @Peter14 @redred413

OpenStudy (anonymous):

@zepdrix @kropot72 @whpalmer4 @msumner @Babyslapmafro

OpenStudy (anonymous):

Let the magnitude of the two forces be A and B. Magnitude of maximum resultant is given as = A+B = 10 Magnitude of minimum resultant is given as = A-B = 6 Solving these two equations ,we get A=8N and B =2N Now we increase A and B by 3N A' = A+3 = 11N B' =B+3 = 5N Now A' and B' are acting perpendicular to each other Therefore resultant is given by \[\sqrt{A'^{2}+B'^{2}}\]

OpenStudy (anonymous):

can you draw ?

OpenStudy (anonymous):

draw what?

OpenStudy (anonymous):

figure for this question..

OpenStudy (dls):

See,we got two forces Question says, The greatest and the least resultant of 2 forces acting at a point is 10Newton and 6Newton respectively. The magnitude of the resultant depends on the angle between the forces since forces are fixed,you can't change them,the variable parameter under your control is theta,i.e the angle. Now the angle here is cos theta, since \[\Huge F_{net}=\sqrt{F_1^2+F_2^2+2F_1F_2\cos \theta}\] You need to know when cos theta is maximum. It is maximum at 0. since cos(0)=1. Magnitude of two forces will be maximum when theta is 0.  and cos of theta is minimum when theta=180 degrees.  Can you try to attempt it now?